OpenStudy (anonymous):

ve for x: 2/5 x + 2 = 10

OpenStudy (diyadiya):

Subtract 2 from both sides and can you tell me what do you get ?

OpenStudy (diyadiya):

2/5+2 -2 =10 -2

OpenStudy (diyadiya):

what is 10 -2 ? @staceyyyy1

OpenStudy (anonymous):

8

OpenStudy (diyadiya):

so (2/5)x =8 Multiply both sides by 5

OpenStudy (diyadiya):

\[\frac{2}{5}x \times 5 = 8 \times 5\]

OpenStudy (anonymous):

8x5=40

OpenStudy (anonymous):

IDK About The Other One.

OpenStudy (diyadiya):